OpenStudy (anonymous):

What number is 144% of 81.25 ?

OpenStudy (anonymous):

please help !!

OpenStudy (anonymous):

144%(81.25)=(144/100)(81.25)=1.44(81.25)=117

OpenStudy (anonymous):

thank u!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

144 is 15% of what number? the answer would be 960??

OpenStudy (anonymous):

Let this number be x, then: 0.15x=144 ---> x=144/0.15=960